# Weak coloring numbers of minor-closed graph classes

### v1

@@ 7.1. Preliminaries
    for every $u\in V(G)$, $T[\{x\in V(T)\mid u\in W_{x}\}]$ is a connected nonnull subtree of $T$, and
@@
For every graph $G$, for every tree decomposition $\mathcal{D}$ of $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, for every positive integer $d$, either
@@
We will need the following technical statement. For a given set $S$, we say that a collection $S_{1},\dots,S_{k}\subseteq S$ is a covering of $S$ if $S_{1}\cup\dots\cup S_{k}=S$.
@@
Let $k$ be a positive integer and let $X$ be a graph. There exists a graph $X^{\prime}$ such that $\operatorname{rtd}_{2}(X^{\prime})\leqslant\operatorname{rtd}_{2}(X)$ and for every sets $S_{1},\dots,S_{k}$ of vertices of $X^{\prime}$ whose union is $V(X^{\prime})$, there exists $i\in[k]$ such that $X^{\prime}$ contains an $\{H\subseteq X^{\prime}\mid\text{$H$ connected and $S_{i}\cap V(H)\neq\emptyset$}\}$-rich model of $X$.
@@
In the case of $h=1$.Let $X^{\prime}=(k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y)$ and $d^{\prime}=1$. In particular, $T_{1,d}(X^{\prime})=K_{1}\oplus((k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y))$. By (iv) and Observation 27
@@
Note that $T_{1,1}(X^{\prime})=K_{1}\oplus X^{\prime}=K_{1}\oplus((k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y))$. Denote by $u$ the vertex of $K_{1}$ in $T_{1,1}(X^{\prime})$ and by $H_{1},\dots,H_{k+1}$ the copies of $T_{h_{1},d_{1}^{\prime}}(Y)$ in $T_{1,1}(X^{\prime})$. Next, let $S_{1},\dots,S_{k}$ be a covering of $T_{1,1}(X^{\prime})$. For every $j\in[k+1]$, there exists $i_{j}\in[k]$ such that $H_{j}$ contains an $\{H\subseteq H_{j}\mid\text{$H$ connected and $S_{i_{j}}\cap V(H)\neq\emptyset$}\}$-rich model $\big{(}B_{j,x}\mid x\in V(X)\big{)}$ of $X$. By the pigeonhole principle, there exist distinct $j_{1},j_{2}\in[k+1]$ such that $i_{j_{1}}=i_{j_{2}}$. Adding a branch set $\{u\}\cup\bigcup_{x\in V(X)}B_{j_{2},x}$ to the model $\big{(}B_{j,x}\mid x\in V(X)\big{)}$ gives a model of $T_{1,d}(X)=K_{1}\oplus X$ in $T_{1,1}(X^{\prime})$. The new branch set contains the root of $T_{1,1}(X^{\prime})$ and corresponds to the root of $T_{1,d}(X)$. Finally, the obtained model is an $\{H\subseteq T_{1,1}(X^{\prime})\mid\text{$H$ connected and $S_{i_{1}}\cap V(H)\neq\emptyset$}\}$-rich model.
Next, suppose that $h>1$ and that the result holds for $h-1$. By induction hypothesis applied to $h-1$, $d$, and $X$, there exists a positive integer $d_{0}^{\prime}$ and a graph $X^{\prime}_{0}$ with $\operatorname{rtd}_{2}(X^{\prime}_{0})\leqslant\operatorname{rtd}_{2}(X)$ such that for each covering $S_{1},\dots,S_{k}$ of $V(T_{h-1,d^{\prime}}(X^{\prime}_{0}))$, there exists $i\in[k]$ such that $T_{h-1,d^{\prime}_{0}}(X^{\prime}_{0})$ contains an $\{H\subseteq T_{h-1,d^{\prime}_{0}}(X^{\prime}_{0})\mid\text{$H$ connected and$S_{i}\cap V(H)\neq\emptyset$}\}$-rich model of $T_{h-1,d}(X)$ whose branch set corresponding to the root of $T_{h-1,d}(X)$ contains the root of $T_{h-1,d^{\prime}}(X^{\prime}_{0})$. Let $d^{\prime}=d^{\prime}_{0}+(dk+1)$ and $X^{\prime}=X^{\prime}_{0}\sqcup\big{(}(k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y)\big{)}$. We claim that $X^{\prime},d^{\prime}$ satisfy the conclusion of the lemma. By (iv) and Observation 27, $\operatorname{rtd}_{2}(X^{\prime})\leqslant\operatorname{rtd}_{2}(X)$.
@@
Let $H_{1},\dots,H_{k+1}$ be the copies of $T_{h_{1},d_{1}^{\prime}}(Y)$ in $X^{\prime}$. For every $j\in[k+1]$ and for every $\ell\in[k]$, let $S_{j,\ell}=\{x\in V(H_{i})\mid i_{x}=\ell\}$. For every $j\in[k+1]$, $S_{j,1},\dots,S_{j,k}$ is a covering of $V(H_{j})$, therefore, there exists $\ell_{j}\in[k]$ such that $H_{j}$ contains an $\{H\subseteq H_{i}\mid\text{$H$ connected and $S_{j,\ell_{j}}\cap V(H)\neq\emptyset$}\}$-rich model $\mathcal{M}_{j}$ of $X$. By pigeonhole principle, there exists distinct $j_{1},j_{2}\in[k+1]$ and $i\in[k]$ such that $\ell_{j_{1}}=\ell_{j_{2}}=i$ for some $i\in[k]$. Let $u$ be the vertex of $K_{1}$ in $k_{1}\oplus X^{\prime}$. Adding a branch set $\{u\}\cup\bigcup\mathcal{M}_{2}$ to the model $\mathcal{M}_{1}$ gives a model of an $\{H\subseteq K_{1}\oplus X^{\prime}\mid\text{$H$ connected and there is $x\in V(H)$ with $i_{x}=i$}\}$-rich model $\big{(}A_{y}\mid y\in V(K_{1}\oplus X)\big{)}$ of $K_{1}\oplus X$ in $K_{1}\oplus X^{\prime}$. For every $y\in V(K_{1}\oplus X$, let $x_{y}\in A_{y}$ be such that $i_{x_{y}}=i$ and let $B_{y}=A_{y}\cup\bigcup\mathcal{M}_{j_{d+1},x_{y}}$. Finally, consider the model obtained from $\big{(}B_{y}\mid y\in V(K_{1}\oplus X)\big{)}$ by adding all the models $\mathcal{M}_{j_{1,x_{y}},x_{y}},\dots,\mathcal{M}_{j_{d,x_{y}},x_{y}}$ for every $y\in V(K_{1}\oplus X)$. We obtain an $\{H\subseteq T_{h,d^{\prime}}(X^{\prime})\mid\text{$H$ connected and $S_{i}\cap V(H)\neq\emptyset$}\}$-rich model of $T_{h,d}(X)$ whose branch set of the root of $T_{h,d}(X)$ contains the root of $T_{h,d^{\prime}}(X^{\prime})$. ∎
@@ 7.2. The base case
This first lemma is a modification of a proof in [3].
@@
Let $d$ be a positive integer. Let $G$ be a connected graph, let $\mathcal{D}=\big{(}T,(W_{x}\mid x\in V(T))\big{)}$ be a tree decomposition of $G$, and let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$ such that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. For every $u\in V(G)$, there is a set $S\subseteq V(G)$ and a layering $(P_{0},\dots,P_{\ell})$ of $G[S]$ with $P_{0}=\{u\}$ such that
@@
We illustrate some parts of the proof in Figure 17. We proceed by induction on $|V(G)|$. If $\mathcal{F}|_{G-\{u\}}$ is empty, then it suffices to take $\ell=0$ and $P_{0}=\{u\}$. In particular, this is the case for $|V(G)|=1$. Therefore, assume $|V(G)|>1$ and $\mathcal{F}|_{G-\{u\}}\neq\emptyset$. Let $\mathcal{F}_{0}$ be the family of all the nonnull connected subgraphs $A$ of $G-\{u\}$ such that $A$ contains some member of $\mathcal{F}$ and $V(A)\cap N_{G}(u)\neq\emptyset$. Since $\mathcal{F}|_{G-\{u\}}\neq\emptyset$ and $G$ is connected, $\mathcal{F}_{0}$ is nonempty.
@@
Let $i,j\in\{0,\dots,\ell\}$ with $i<j$ and assume that there is an edge a vertex in $P_{i}$ and a vertex in $P_{j}$ in $G$. If $i\geqslant 2$, then $P_{i}\subseteq P^{\prime}_{i-1}$ and $P_{j}=P^{\prime}_{j-1}$, which implies $|i-j|\leqslant 1$ since $(P_{0}^{\prime},\dots,P_{\ell^{\prime}}^{\prime})$ is a layering of $G^{\prime}[S^{\prime}]$. Otherwise, $i\in\{0,1\}$. If $i=0$, then $j=1$ since $u$ has no neighbors in $W$. If $i=1$, then $j=2$ since $N_{G^{\prime}}(u^{\prime})\subseteq P_{1}^{\prime}$. It follows that $(P_{0},\dots,P_{\ell})$ is a layering of $G[S]$.
Let $F\in\mathcal{F}$. If $V(F)\cap(\{u\}\cup Z)\neq\emptyset$, then $V(F)\cap S\neq\emptyset$. Otherwise, $F\subseteq C$ for some component of $G-(\{u\}\cup Z)$, and in particular, $C\notin\mathcal{F}^{\prime}$. In this case, $N_{G}(u)\cap V(C)$, hence, $C\in\mathcal{C}_{0}$, thus, $F\in\mathcal{F}|_{G[W]}$, and finally, $V(F)\cap S^{\prime}\neq\emptyset$ by (a’). This proves (a).
@@
Recall that in the base case of the proof of Theorem 5 (see Section 5), we applied the ideas required to prove that positive integer $r$ and for every path $P$, $\operatorname{wcol}_{r}(P)\leqslant 1+\lceil\log r\rceil$. We illustrated these ideas in Figure 10. In this section, it suffices to use this result as a black box. We state it here for reference.
@@
For every positive integer $r$ and for every path $P$, $\operatorname{wcol}_{r}(P)\leqslant 1+\lceil\log r\rceil$.
@@
Let $k,d$ be positive integers. For every graph $G$, for every tree decomposition $\mathcal{D}$ of $G$ of width at most $k-1$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$, then there is a set $S\subseteq V(G)$ such that
@@
    for every component $C$ of $G-S$, $N_{G}(V(C))$ is contained in the union of at most $d$ bags of $\mathcal{D}$;
@@
Let $G$ be a graph, let $\mathcal{D}$ be a tree decomposition of $G$ of width at most $k-1$, let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$, and suppose that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. Let $r$ be an integer with $r\geqslant 2$. By considering the components of $G$ independently, we can assume that $G$ is connected. Let $u$ be an arbitrary vertex in $G$. Lemma 40 applied to $G$, $\mathcal{D}$, $\mathcal{F}$, and $u$ gives $S\subseteq V(G)$, a layering $(P_{0},\dots,P_{\ell})$ of $G[S]$ with $P_{0}=\{u\}$ such that
@@
Note that (a) holds by 40.(a) and (b) holds by 40.(b). In order to conclude the proof, it suffices to show (c).
@@
Let $k,h,d$ be positive integers with $h\geqslant 2$. There is an integer $\bar{c}_{0}(h,d)$ such that for every graph $G$, for every tree decomposition $\mathcal{D}$ of $G$ of width at most $k-1$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$, then there is a set $S\subseteq V(G)$ such that
@@
    for every component $C$ of $G-S$, $N_{G}(V(C))$ is contained in the union of at most $d(h-1)+\binom{h-1}{2}$ bags of $\mathcal{D}$;
@@
We proceed by induction on $h$. For $h=2$, the result holds by Lemma 42 setting $\bar{c}_{0}(h,d)=6dk$, since $d(2-1)+\binom{2-1}{2}=d$. Next, assume $h\geqslant 3$ and that $\bar{c}(h,d)$ witnesses the assertion for $h-1$. Let $\bar{c}_{0}(h,d)=6d+\bar{c}_{0}(h-1,d+1)$.
Let $G$ be a graph and let $\mathcal{D}$ be a tree decomposition of $G$ of width at most $k-1$. Let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$. Let $r$ be an integer with $r\geqslant 2$. Let $\mathcal{F}^{\prime}$ be the family of all the connected subgraphs $H$ of $G$ such that $H$ contains an $\mathcal{F}|_{H}$-rich model of $F_{h-1,d+1}$. We claim that there is no $\mathcal{F}^{\prime}$-rich model of $F_{2,d}$ in $G$. Suppose to the contrary that $\big{(}B_{x}\mid x\in V(F_{2,d})\big{)}$ is such a model. Let $s$ be the root of $F_{2,d}$ and let $s^{\prime}$ be the root of $F_{h-1,d}$. For every $x\in V(F_{2,d})-\{s\}$, by Lemma 32, there is an $\mathcal{F}$-rich model $\big{(}C_{y}\mid y\in V(F_{h-1,d})\big{)}$ of $F_{h-1,d}$ in $G[B_{x}]$ such that $C_{s^{\prime}}$ contains a vertex of $N_{G}(B_{s})\cap B_{x}$. The union of these models together with $B_{s}$ yields an $\mathcal{F}$-rich model of $F_{h,d}$ in $G$, which is a contradiction. Note that this is exactly the same argument as in Lemma 33, see Figure 11.
@@
    for every component $C$ of $G-S_{0}$, $N_{G}(V(C))$ is contained in the union of at most $d$ bags of $\mathcal{D}$;
@@
    for every component $C^{\prime}$ of $C-S_{C}$, $N_{C}(V(C^{\prime}))$ is contained in the union of at most $(d+1)(h-2)+\binom{h-2}{2}=d(h-1)+\binom{h-1}{2}-d$ bags of $\mathcal{D}_{C}$;
@@
We claim that (a)-(c) hold. Let $F\in\mathcal{F}$. If $V(F)\cap S_{0}=\emptyset$, then $V(F)\subseteq V(C)$ for some component $C$ of $G-S_{0}$. In particular, $F\in\mathcal{F}|_{C}$, and thus, by (a’), $V(F)\cap S_{C}\neq\emptyset$, which proves (a). For every component $C^{\prime}$ of $G-S$, there exists a component $C$ of $G-S_{0}$ such that $C^{\prime}\subseteq C$ and by 42.(b) and (b’), $N_{G}(C^{\prime})$ is contained in at most $d+\left(d(h-1)+\binom{h-1}{2}-d\right)=d(h-1)+\binom{h-1}{2}$ bags of $\mathcal{D}$, which implies (b).
@@
We now show that using Lemma 37, the constant $d(h-1)+\binom{h-1}{2}$ in (b) in Lemma 43 can be reduced to two.
@@
Let $k,h,d$ be positive integers with $h\geqslant 2$. For every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$, then there is a set $S\subseteq V(G)$ such that
@@
    $\operatorname{wcol}_{r}(G,S)\leqslant\left(\bar{c}_{0}(h,d)+2dh^{2}\right)k\cdot\log r$ for every integer $r$ with $r\geqslant 2$, where $\bar{c}_{0}(h,d)$ is the constant from Lemma 43.
@@
Let $G$ be a graph and let $\mathcal{D}$ be a tree decomposition of $G$ of width at most $k-1$. By Lemma 36, we can assume that $\mathcal{D}$ is a natural tree decomposition of $G$. By Lemma 42 applies to $G$, $\mathcal{D}$, and $\mathcal{F}$, there is a set $S_{0}\subseteq V(G)$ such that
@@
    for every component $C$ of $G-S_{0}$, $N_{G}(V(C))$ is contained in the union of at most $d(h-1)+\binom{h-1}{2}$ bags of $\mathcal{D}$;

### v2

@@ 7.1. Preliminaries
    for every $u\in V(G)$, $T[\{x\in V(T)\mid u\in W_{x}\}]$ is a connected subtree of $T$, and
@@
For every graph $G$, for every tree decomposition $\mathcal{D}$ of $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, for every positive integer $d$, either
@@
We need the following technical statement. For a given set $S$, we say that a collection $S_{1},\dots,S_{k}\subseteq S$ is a covering of $S$ if $S_{1}\cup\dots\cup S_{k}=S$.
@@
Let $k$ be a positive integer and let $X$ be a graph. There exists a graph $X^{\prime}$ such that $\operatorname{rtd}_{2}(X^{\prime})\leqslant\operatorname{rtd}_{2}(X)$ and for each covering $S_{1},\dots,S_{k}$ of $V(X^{\prime})$, there exists $i\in[k]$ such that $X^{\prime}$ contains an $\{H\subseteq X^{\prime}\mid\text{$H$ connected and $S_{i}\cap V(H)\neq\emptyset$}\}$-rich model of $X$.
@@
In the case of $h=1$. Let $X^{\prime}=(k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y)$ and $d^{\prime}=1$. In particular, $T_{1,d}(X^{\prime})=K_{1}\oplus((k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y))$. By (iv) and Lemma 27
@@
Note that $T_{1,1}(X^{\prime})=K_{1}\oplus X^{\prime}=K_{1}\oplus((k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y))$. Denote by $u$ the vertex of $K_{1}$ in $T_{1,1}(X^{\prime})$ and by $H_{1},\dots,H_{k+1}$ the copies of $T_{h_{1},d_{1}^{\prime}}(Y)$ in $T_{1,1}(X^{\prime})$. Next, let $S_{1},\dots,S_{k}$ be a covering of $T_{1,1}(X^{\prime})$. For every $j\in[k+1]$, there exists $i_{j}\in[k]$ such that $H_{j}$ contains an $\{H\subseteq H_{j}\mid\text{$H$ connected and $S_{i_{j}}\cap V(H)\neq\emptyset$}\}$-rich model $\big{(}B_{j,x}\mid x\in V(X)\big{)}$ of $X$. By the pigeonhole principle, there exist distinct $j_{1},j_{2}\in[k+1]$ such that $i_{j_{1}}=i_{j_{2}}$. Let $i=i_{j_{1}}=i_{j_{2}}$. Adding a branch set $\{u\}\cup\bigcup_{x\in V(X)}B_{j_{2},x}$ to the model $\big{(}B_{j_{1},x}\mid x\in V(X)\big{)}$ gives a model of $T_{1,d}(X)=K_{1}\oplus X$ in $T_{1,1}(X^{\prime})$. The new branch set contains the root of $T_{1,1}(X^{\prime})$ and corresponds to the root of $T_{1,d}(X)$. Finally, the obtained model is an $\{H\subseteq T_{1,1}(X^{\prime})\mid\text{$H$ connected and $S_{i}\cap V(H)\neq\emptyset$}\}$-rich model of $T_{1,d}(X)$.
Next, suppose that $h>1$ and that the result holds for $h-1$. By induction hypothesis applied to $h-1$, $d$, and $X$, there exists a positive integer $d_{0}^{\prime}$ and a graph $X^{\prime}_{0}$ with $\operatorname{rtd}_{2}(X^{\prime}_{0})\leqslant\operatorname{rtd}_{2}(X)$ such that for each covering $S_{1},\dots,S_{k}$ of $V(T_{h-1,d^{\prime}}(X^{\prime}_{0}))$, there exists $i\in[k]$ such that $T_{h-1,d^{\prime}_{0}}(X^{\prime}_{0})$ contains an $\{H\subseteq T_{h-1,d^{\prime}_{0}}(X^{\prime}_{0})\mid\text{$H$ connected and$S_{i}\cap V(H)\neq\emptyset$}\}$-rich model of $T_{h-1,d}(X)$ whose branch set corresponding to the root of $T_{h-1,d}(X)$ contains the root of $T_{h-1,d^{\prime}}(X^{\prime}_{0})$. Let $d^{\prime}=d^{\prime}_{0}+(dk+1)$ and $X^{\prime}=X^{\prime}_{0}\sqcup\big{(}(k+1)\cdot T_{h_{1},d_{1}^{\prime}}(Y)\big{)}$. We claim that $X^{\prime},d^{\prime}$ satisfy the conclusion of the lemma. By (iv) and Lemma 27, $\operatorname{rtd}_{2}(X^{\prime})\leqslant\operatorname{rtd}_{2}(X)$.
@@
Let $H_{1},\dots,H_{k+1}$ be the copies of $T_{h_{1},d_{1}^{\prime}}(Y)$ in $X^{\prime}$. For every $j\in[k+1]$ and for every $\ell\in[k]$, let $S_{j,\ell}=\{x\in V(H_{j})\mid i_{x}=\ell\}$. For every $j\in[k+1]$, $S_{j,1},\dots,S_{j,k}$ is a covering of $V(H_{j})$, therefore, there exists $\ell_{j}\in[k]$ such that $H_{j}$ contains an $\{H\subseteq H_{i}\mid\text{$H$ connected and $S_{j,\ell_{j}}\cap V(H)\neq\emptyset$}\}$-rich model $\mathcal{M}_{j}$ of $X$. By pigeonhole principle, there exist distinct $j_{1},j_{2}\in[k+1]$ and $i\in[k]$ such that $\ell_{j_{1}}=\ell_{j_{2}}$. Let $i=\ell_{j_{1}}=\ell_{j_{2}}$. Let $u$ be the vertex of $K_{1}$ in $k_{1}\oplus X^{\prime}$. Adding a branch set $\{u\}\cup\bigcup\mathcal{M}_{j_{2}}$ to the model $\mathcal{M}_{j_{1}}$ gives a model of an $\{H\subseteq K_{1}\oplus X^{\prime}\mid\text{$H$ connected and there is $x\in V(H)$ with $i_{x}=i$}\}$-rich model $\big{(}A_{y}\mid y\in V(K_{1}\oplus X)\big{)}$ of $K_{1}\oplus X$ in $K_{1}\oplus X^{\prime}$. For every $y\in V(K_{1}\oplus X$, let $x_{y}\in A_{y}$ be such that $i_{x_{y}}=i$ and let $B_{y}=A_{y}\cup\bigcup\mathcal{M}_{j_{d+1},x_{y}}$. Finally, consider the model obtained from $\big{(}B_{y}\mid y\in V(K_{1}\oplus X)\big{)}$ by adding all the models $\mathcal{M}_{j_{1,x_{y}},x_{y}},\dots,\mathcal{M}_{j_{d,x_{y}},x_{y}}$ for every $y\in V(K_{1}\oplus X)$. We obtain an $\{H\subseteq T_{h,d^{\prime}}(X^{\prime})\mid\text{$H$ connected and $S_{i}\cap V(H)\neq\emptyset$}\}$-rich model of $T_{h,d}(X)$ whose branch set of the root of $T_{h,d}(X)$ contains the root of $T_{h,d^{\prime}}(X^{\prime})$. ∎
@@ 7.2. The base case
This first lemma is a modification of a proof in [3, Lemma 8].
@@
Let $d$ be a positive integer. Let $G$ be a connected graph, let $\mathcal{D}=\big{(}T,(W_{x}\mid x\in V(T))\big{)}$ be a tree decomposition of $G$, and let $\mathcal{F}$ be a family of connected subgraphs of $G$ such that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. For every $u\in V(G)$, there is a set $S\subseteq V(G)$ and a layering $(P_{0},\dots,P_{\ell})$ of $G[S]$ with $P_{0}=\{u\}$ such that
@@
We illustrate some parts of the proof in Figure 17. We proceed by induction on $|V(G)|$. If $\mathcal{F}|_{G-\{u\}}$ is empty, then it suffices to take $\ell=0$ and $P_{0}=\{u\}$. In particular, this is the case for $|V(G)|=1$. Therefore, assume $|V(G)|>1$ and $\mathcal{F}|_{G-\{u\}}\neq\emptyset$. Let $\mathcal{F}_{0}$ be the family of all the connected subgraphs $A$ of $G-\{u\}$ such that $A$ contains some member of $\mathcal{F}$ and $V(A)\cap N_{G}(u)\neq\emptyset$. Since $\mathcal{F}|_{G-\{u\}}\neq\emptyset$ and $G$ is connected, $\mathcal{F}_{0}$ is nonempty.
@@
Let $i,j\in\{0,\dots,\ell\}$ with $i<j$ and assume that there is an edge incident to a vertex in $P_{i}$ and a vertex in $P_{j}$ in $G$. If $i\geqslant 2$, then $P_{i}\subseteq P^{\prime}_{i-1}$ and $P_{j}=P^{\prime}_{j-1}$, which implies $|i-j|\leqslant 1$ since $(P_{0}^{\prime},\dots,P_{\ell^{\prime}}^{\prime})$ is a layering of $G^{\prime}[S^{\prime}]$. Otherwise, $i\in\{0,1\}$. If $i=0$, then $j=1$ since $u$ has no neighbors in $W$. If $i=1$, then $j=2$ since $N_{G^{\prime}}(u^{\prime})\subseteq P_{1}^{\prime}$. It follows that $(P_{0},\dots,P_{\ell})$ is a layering of $G[S]$.
Let $F\in\mathcal{F}$. If $V(F)\cap(\{u\}\cup Z)\neq\emptyset$, then $V(F)\cap S\neq\emptyset$. Otherwise, $F\subseteq C$ for some component of $G-(\{u\}\cup Z)$, and in particular, $C\notin\mathcal{F}_{0}$. In this case, $N_{G}(u)\cap V(C)$, hence, $C\in\mathcal{C}_{0}$, thus, $F\in\mathcal{F}|_{G[W]}$, and finally, $V(F)\cap S^{\prime}\neq\emptyset$ by (a’). This proves (a).
@@
Recall that in the base case of the proof of Theorem 5 (see Section 5), we applied the ideas required to prove that positive integer $r$ and for every path $P$, $\operatorname{wcol}_{r}(P)\leqslant 2+\lceil\log r\rceil$. We illustrated these ideas in Figure 10. In this section, it suffices to use this result as a black box. We state it here for reference.
@@
For every positive integer $r$ and for every path $P$, $\operatorname{wcol}_{r}(P)\leqslant 2+\lceil\log r\rceil$.
@@
Let $k,d$ be positive integers. For every graph $G$, for every tree decomposition $\mathcal{D}$ of $G$ of width at most $k-1$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$, then there is a set $S\subseteq V(G)$ such that
@@
    for every component $C$ of $G-S$, $N_{G}(V(C))$ is contained in the union of at most $2d$ bags of $\mathcal{D}$;
@@
Let $G$ be a graph, let $\mathcal{D}$ be a tree decomposition of $G$ of width at most $k-1$, let $\mathcal{F}$ be a family of connected subgraphs of $G$, and suppose that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. Let $r$ be an integer with $r\geqslant 2$. By considering the components of $G$ independently, we can assume that $G$ is connected. Let $u$ be an arbitrary vertex in $G$. Lemma 40 applied to $G$, $\mathcal{D}$, $\mathcal{F}$, and $u$ gives $S\subseteq V(G)$, a layering $(P_{0},\dots,P_{\ell})$ of $G[S]$ with $P_{0}=\{u\}$ such that
@@
Note that (a) holds by 40.(a) and (b) holds by 40.(b) and 40.(c). In order to conclude the proof, it suffices to show (c).
@@
Let $k,h,d$ be positive integers with $h\geqslant 2$. There is an integer $\bar{c}_{0}(h,d)$ such that for every graph $G$, for every tree decomposition $\mathcal{D}$ of $G$ of width at most $k-1$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$, then there is a set $S\subseteq V(G)$ such that
@@
    for every component $C$ of $G-S$, $N_{G}(V(C))$ is contained in the union of at most $2d(h-1)+2\binom{h-1}{2}$ bags of $\mathcal{D}$;
@@
We proceed by induction on $h$. For $h=2$, the result holds by Lemma 42 setting $\bar{c}_{0}(h,d)=6dk$, since $2d(2-1)+2\binom{2-1}{2}=2d$. Next, assume $h\geqslant 3$ and that $\bar{c}_{0}(h,d)$ witnesses the assertion for $h-1$. Let $\bar{c}_{0}(h,d)=6d+\bar{c}_{0}(h-1,d+1)$.
Let $G$ be a graph and let $\mathcal{D}$ be a tree decomposition of $G$ of width at most $k-1$. Let $\mathcal{F}$ be a family of connected subgraphs of $G$. Let $r$ be an integer with $r\geqslant 2$. Let $\mathcal{F}^{\prime}$ be the family of all the connected subgraphs $H$ of $G$ such that $H$ contains an $\mathcal{F}|_{H}$-rich model of $F_{h-1,d+1}$. We claim that there is no $\mathcal{F}^{\prime}$-rich model of $F_{2,d}$ in $G$. Suppose to the contrary that $\big{(}B_{x}\mid x\in V(F_{2,d})\big{)}$ is such a model. Let $s$ be the root of $F_{2,d}$ and let $s^{\prime}$ be the root of $F_{h-1,d}$. For every $x\in V(F_{2,d})-\{s\}$, by Lemma 32, there is an $\mathcal{F}$-rich model $\big{(}C_{y}\mid y\in V(F_{h-1,d})\big{)}$ of $F_{h-1,d}$ in $G[B_{x}]$ such that $C_{s^{\prime}}$ contains a vertex of $N_{G}(B_{s})\cap B_{x}$. The union of these models together with $B_{s}$ yields an $\mathcal{F}$-rich model of $F_{h,d}$ in $G$, which is a contradiction. Note that this is exactly the same argument as in Lemma 33, see Figure 11.
@@
    for every component $C$ of $G-S_{0}$, $N_{G}(V(C))$ is contained in the union of at most $2d$ bags of $\mathcal{D}$;
@@
    for every component $C^{\prime}$ of $C-S_{C}$, $N_{C}(V(C^{\prime}))$ is contained in the union of at most $2(d+1)(h-2)+2\binom{h-2}{2}=2d(h-1)+2\binom{h-1}{2}-2d$ bags of $\mathcal{D}_{C}$;
@@
We claim that (a)-(c) hold. Let $F\in\mathcal{F}$. If $V(F)\cap S_{0}=\emptyset$, then $V(F)\subseteq V(C)$ for some component $C$ of $G-S_{0}$. In particular, $F\in\mathcal{F}|_{C}$, and thus, by (a’), $V(F)\cap S_{C}\neq\emptyset$, which proves (a). For every component $C^{\prime}$ of $G-S$, there exists a component $C$ of $G-S_{0}$ such that $C^{\prime}\subseteq C$ and by 42.(b) and (b’), $N_{G}(C^{\prime})$ is contained in at most $2d+\left(2d(h-1)+2\binom{h-1}{2}-2d\right)=2d(h-1)+2\binom{h-1}{2}$ bags of $\mathcal{D}$, which implies (b).
@@
We now show that using Lemma 37, the constant $2d(h-1)+2\binom{h-1}{2}$ in (b) in Lemma 43 can be reduced to two.
@@
Let $k,h,d$ be positive integers with $h\geqslant 2$. For every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$, then there is a set $S\subseteq V(G)$ such that
@@
    $\operatorname{wcol}_{r}(G,S)\leqslant\left(\bar{c}_{0}(h,d)+4dh^{2}\right)k\cdot\log r$ for every integer $r$ with $r\geqslant 2$, where $\bar{c}_{0}(h,d)$ is the constant from Lemma 43.
@@
Let $G$ be a graph and let $\mathcal{D}$ be a tree decomposition of $G$ of width at most $k-1$. We may assume $G$ is connected. By Lemma 36, we can assume that $\mathcal{D}$ is a natural tree decomposition of $G$. By Lemma 42 applies to $G$, $\mathcal{D}$, and $\mathcal{F}$, there is a set $S_{0}\subseteq V(G)$ such that
@@
    for every component $C$ of $G-S_{0}$, $N_{G}(V(C))$ is contained in the union of at most $2d(h-1)+2\binom{h-1}{2}$ bags of $\mathcal{D}$;